# Coach not crutch: AI assistance can enhance rather than hinder skill development

### v3

@@ Discussion
Contrary to lay forecasters’ expectations (Studies 1 and 3), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 5), more than those who had access to Google cover letter examples, and even more than those who received personalized feedback from professional human editors (Study 4). Specifically, participants who had practiced with AI, subsequently wrote cover letters (without AI assistance) that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later. Learning gains were not the result of greater effort; in fact, participants who had practiced with AI expended less effort during practice than those who had practiced alone. Instead, these gains were at least in part explained by exposure to high-quality, just-in-time personalized examples: participants who merely viewed an AI-generated cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 5).
Study 1 forecasters who believed AI would harm learning cited reduced effort as the mechanism—and prior research supports this concern (28, 29). All things equal, the more effort a learner invests in a given practice task, the more they learn. So why did practicing with AI improve writing skill more than practicing alone, even while reducing both the quantity and quality of effort? Likewise, why did writers who googled how to compose a cover letter work longer and harder than writers who practiced with AI, yet learn less? We propose a third mechanism that offsets decrements in human effort: the utility of the practice task itself. See Figure 8. Perhaps the excellent personalized examples provided by AI are analogous to demonstrations of an Olympic coach. The higher the utility of the practice task, the faster the learning.
While not immediately apparent to forecasters, teachers (11), and the general population (8), the pedagogical benefits of viewing AI-generated examples are consistent with prior research on learning. Research has shown that humans are especially adept at observing, imitating, and learning from others (30, 31, 32). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (33).
Future Directions
@@
Three promising directions for future research are worth highlighting. First, it is not clear whether the observed benefits of AI examples would be as large in domains other than professional writing. At a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting). In other domains, however, merely observing a solution may be less informative. For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it.
Second, additional research is needed to explore moderators of learning from AI tools. Certain metacognitive strategies may enhance the benefits of interacting with AI. For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21). Similarly, correlational evidence suggests that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22). Conversely, other factors may minimize learning benefits. In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI. During practice, therefore, participants were incentivized to prioritize gains in acquired skill over performance in the moment. In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI. More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without increasing task utility.
@@
Third, in our experimental paradigm, participants interacted with the AI tool only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (34). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI tools.
Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although the AI writing tool reduced the effort users invested in practicing, it nevertheless accelerated their skill development. The underappreciated efficacy of timely and tailored examples carries practical implications: Many AI tools designed to support learning are explicitly programmed not to “give away” answers. However, it may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
Decades before the advent of generative AI, the legendary UCLA basketball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (35). Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain. In doing so, AI has the potential not only to boost productivity but also to democratize opportunities for building human capital at scale.
@@ Methods
Study 1 was preregistered (#191800). The moderation analyses were not pre-registered.
Study 2 was pre-registered (#201140). The pairwise comparisons analysis was pre-registered separately (#205316). After preregistration, we considered a keystrokes-per-minute metric of effort intensity rather than duration; thus, this analysis was not preregistered.
Study 3 was pre-registered (#236487). Analyses on predicted effectiveness of AI vs. human feedback and the mediation analysis were not preregistered.
Study 4 was pre-registered (#239157). The pairwise comparisons analysis was pre-registered separately (#243115). The keystrokes-per-minute metric of effort was not pre-registered.
Study 5 was pre-registered (#197704). The one-day follow-up was collected in three batches: a pilot, followed by second and third batches. We pre-registered batch 2 (#199451) but report pooled results in the main text, with separate batch details available in Supplementary Information Section F5. The pairwise comparisons analysis was pre-registered separately (#205315). The keystrokes-per-minute metric of effort was not pre-registered.
@@
We sampled participants from Prolific from all our studies. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
Participants in Study 1 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2). They were predominantly white (75%). A small proportion were students (13%), and most were employed (62%).
In Study 2, the sample was more evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5). Over half of the sample (58%) was White, with the rest comprising Black (33%), Latino (6%), and Asian (5%) participants. Most participants (77%) had college degrees. Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
Participants in Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4). They were predominantly white (67%). A small proportion were students (19%), and most were employed (56%).
In Study 4, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8). Over half of the sample (74%) was White, with the rest comprising Black (16%), Latino (8%), and Asian (7%) participants. Most participants (74%) had college degrees. Most participants (94%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants: 17% had never used AI assistants, 40% had tried them but hardly ever used them, and 43% used them at least weekly.
Study 5 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6). Over half of the sample (64%) was White, with the remainder comprising Black (24%), Latino (8%), and Asian (6%) participants. Most participants (74%) had college degrees. Most participants (91%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (40% had tried them, but hardly ever used them, 42% used them at least a few times per week, and 18% had never used AI assistants before).
@@
All studies followed the same general structure with study-specific variations. Participants first saw an introductory screen and completed a brief questionnaire reporting demographics, prior experience with AI writing tools, motivation to improve their writing, and perceived writing skill. They then completed a short pre-test (2 minutes) in which they edited a poorly written cover letter, followed by a brief lesson on the five principles of effective writing.
In Studies 2 and 5, participants were then randomized to a practice condition (or skipped practice in the no-practice control). During practice, participants rewrote a new cover letter or, in the example-only condition, viewed an AI-rewritten version of the letter. This example was not explicitly labeled as AI-generated. After practice, participants saw either their own rewritten letter or the AI-generated example, along with AI-generated feedback highlighting one suggested improvement (see Supplementary Information, Section A3). Immediately after this feedback, participants rated how much they had learned and how hard they had worked so far. They then completed a 7-minute incentivized test of writing skill by editing a new cover letter without access to AI tools. To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality. Finally, participants could request optional feedback and were asked whether they had used outside resources during the test. A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S6, S18, and S27).
In Study 4, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search. Participants completed the same baseline questionnaire, pretest, and writing lesson. They then edited a new cover letter without any assistance and were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional editors, (3) practice with a Google Search for cover letter examples and tips. In the second condition condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter. The following day, participants were asked to revise their original cover letter using the resources from their assigned condition. All participants then completed a final writing test by rewriting a new cover letter without assistance, followed by an exit questionnaire.
@@
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (36, 37). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Tables S6, S18, and S27.
To validate these ratings, the first author and a trained research assistant took a sample of n = 100 cover letters from Study 2, and rated them on the 5 principles. The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001). These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 4, 5, respectively.)
To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters. We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6, S18, and S27), suggesting that our effects are not explainable by same-model bias.
Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 4 and 5, respectively).
@@
As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness). We used logistic regression to predict whether participants chose to see optional test feedback from condition, controlling for pretest score and baseline characteristics. Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1. When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (38).
@@ E Results Study 4
- E2 Details on professional editors
- E3 Details on participants’ search behavior
- E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
- E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
- E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
- E7 Results were not moderated by individual differences
- E8 Pairwise comparisons
F Results Study 5
- F1 Randomization, Balance, and Missingness
- F2 AI examples improve writing skill
- F3 Seeing AI examples was less effortful
- F4 Seeing an AI example did not discourage motivation for future learning
- F5 The benefits of seeing an AI example were just as large a day later
- F6 Seeing AI examples was equally effective across subgroups
- F7 Pairwise Comparisons
@@ Appendix A Additional methods
We used a randomly selected sample of 100 cover letters from Study 2 to validate the AI ratings. Two trained raters independently rated each of the 5 principles. The inter-rater correlation was r = .74, p < .001. The human raters correlated with the AI-generated ratings satisfactorily (r = .70, p < .001). See Figure S1.
@@ Appendix B Results Study 1
| Predictor | OR | SE | z | p-value |
| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
| Employment status | ||||
| Other | 0.770 | 0.687 | -0.382 | 0.703 |
| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |

### v4

@@ Discussion
Although many believe that using AI undermines effort and learning, we found in our experiments that participants assigned to practice writing cover letters with AI learned more than those assigned to practice on their own. Access to AI also led to more learning than did access to Google search for cover letter tips and examples, or even personalized feedback from professional human editors.
Lay beliefs about AI and effort were more accurate. Participants given access to an AI tool expended less effort while practicing than those who practiced alone or who practiced with access to Google tips and examples. Not surprisingly, practicing with AI was about as effortful as practicing with feedback from a professional human editor.
What accounts for the learning gains from practicing with AI? We argue that AI improves the learning environment by providing high-quality, just-in-time, personalized examples that illustrate otherwise abstract principles. As shown in Figure 1, even if AI reduces the effort people expend during practice, it may nevertheless improve learning by enhancing the learning environment. This mechanism is consistent with a large literature showing that humans are especially adept at learning by observation (bandura1971, meltzoff2005imitation, lyons2007hidden). Relatedly, individualized tutoring is a much more effective learning environment than whole group instruction (bloom_2_1984)—and experts in diverse domains engage in deliberate practice, acquiring increasingly more sophisticated mental representations provided by their teachers and coaches (anders2008deliberate). Compared to textbooks, conventional computer-tutoring programs (e.g., (anderson1995cognitive)), and even human instructors, AI systems are especially adept at instantly generating specific, high-quality examples tailored to a learner’s momentary needs. Indeed, we found that AI-generated examples of cover letters during the practice phase were better written and more likely to elicit interviews than cover letters produced by professional writers and editors.
Three promising directions for future research are worth highlighting.
@@
First, it is not clear whether the observed benefits of AI examples would generalize to other domains like math (kumar2023, bastani2024) and computer programming (nie2024gpt). At a glance, a single AI-generated example illustrates the principles of effective professional writing. In other domains, merely observing an example may be less informative. For instance, the final answer to a math or computer programming problem does not often reveal the process that produced it.
Second, additional research is needed to explore moderators. There may be metacognitive strategies that enhance the benefits of writing with AI. For instance, trying on one’s own before using AI has been shown to improve learning more than the converse (cf. (kumar2023)). Conversely, time pressure and competing priorities around performance and learning may minimize the learning benefits. More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without improving the learning experience.
@@
Third, in our experimental paradigm, participants interacted with AI only once. It is common, however, to use AI repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (shin2023). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI.
Taken together, our findings challenge the widely held view that productivity boosts from AI inevitably hamper the development of human skill. By illustrating otherwise abstract principles with examples personalized to the precise needs of the learner at the moment, AI can, at least in some use cases, improve the quality of the learning environment. Rather than viewing AI as a universal threat to human capital, our results provide an existence proof that AI can support the development of skills at scale.
@@ Methods
Study 1 was a descriptive survey and was not preregistered.
Study 2 was pre-registered (#201140). The pairwise comparisons analysis was pre-registered separately (#205316). After preregistration, we considered a keystrokes-per-minute metric of effort intensity rather than duration; thus, this analysis was not preregistered. The forecasting data was preregistered (#191800), but not the moderation analysis within it.
Study 3 was pre-registered (#239157). The pairwise comparisons analysis (#243115) and the comparison between editor rewrites and counterfactual AI rewrites (#255792) were pre-registered separately. The keystrokes-per-minute metric of effort was not pre-registered. The willingness to pay data was preregistered (#236487), but not the analyses on the predicted effectiveness of AI vs. human feedback and the mediation analysis.
Study 4 was pre-registered (#197704). The one-day follow-up was collected in three batches: a pilot, followed by second and third batches. We pre-registered batch 2 (#199451), but report pooled results in the main text, with separate batch details available in Supplementary Information Section E5. The pairwise comparisons analysis was pre-registered separately (#205315). The keystrokes-per-minute metric of effort was not pre-registered.
See all preregistrations here.
@@
Participants in Study 1 were a nationally representative sample of young adults aged 18-28 (M = 25.3, SD = 2.4) recruited through the Gallup organization. The sample comprised 38% men, 57% women, and 5% nonbinary participants. The majority (64%) were White, with the remainder comprising Hispanic (15%), Asian (10%), Black (10%), and other race (1%) participants. We used sample weights to ensure national representativeness of the data for all analyses.
We sampled participants from Prolific for all our experiments. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
In Study 2, the sample was somewhat evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5). Over half of the sample (58%) was White, with the rest comprising Black (33%), Latino (6%), and Asian (5%) participants. Most participants (77%) had college degrees. Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before). Participants forecasting the results of Study 2 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2). They were predominantly white (75%). A small proportion were students (13%), and most were employed (62%).
In Study 3, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8). Over half of the sample (74%) was White, with the rest comprising Black (16%), Latino (8%), and Asian (7%) participants. Most participants (74%) had college degrees. Most participants (94%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants: 17% had never used AI assistants, 40% had tried them but hardly ever used them, and 43% used them at least weekly. Participants forecasting Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4). They were predominantly white (67%). A small proportion were students (19%), and most were employed (56%).
Study 4 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6). Over half of the sample (64%) was White, with the remainder comprising Black (24%), Latino (8%), and Asian (6%) participants. Most participants (74%) had college degrees. Most participants (91%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (40% had tried them, but hardly ever used them, 42% used them at least a few times per week, and 18% had never used AI assistants before).
@@
In Study 1, participants completed a brief survey about their AI usage in the past month and their general attitudes toward AI and learning. Participants rated their agreement with statements about AI’s impact on learning, intelligence, and work habits using 6-point Likert-type scales (1 = Strongly disagree, 6 = Strongly agree).
All experiments (Studies 2–4) followed the same general structure with study-specific variations. Participants first saw an introductory screen and completed a brief questionnaire reporting demographics, prior experience with AI writing tools, motivation to improve their writing, and perceived writing skill. They then completed a short pre-test (2 minutes) in which they edited a poorly written cover letter, followed by a brief lesson on the five principles of effective writing.
In Studies 2 and 4, participants were then randomized to a practice condition (or skipped practice in the no-practice control). During practice, participants rewrote a new cover letter or, in the example-only condition, viewed an AI-rewritten version of the letter. This example was not explicitly labeled as AI-generated. After practice, participants saw either their own rewritten letter or the AI-generated example, along with AI-generated feedback highlighting one suggested improvement (see Supplementary Information, Section A3). Immediately after this feedback, participants rated how much they had learned and how hard they had worked so far. They then completed a 7-minute incentivized test of writing skill by editing a new cover letter without access to AI tools. To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality. Finally, participants could request optional feedback and were asked whether they had used outside resources during the test. A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S7, S20, and S29).
In Study 3, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search. Participants completed the same baseline questionnaire, pretest, and writing lesson. They then edited a new cover letter without any assistance and were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional editors, (3) practice with a Google Search for cover letter examples and tips. In the second condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter. The following day, participants were asked to revise their original cover letter using the resources from their assigned condition. All participants then completed a final writing test by rewriting a new cover letter without assistance, followed by an exit questionnaire.
@@
For Studies 2-4, as per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (rathje2024, hackl2023). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Tables S7, S20, and S29.
To validate these ratings, the first author and a trained research assistant took a sample of n = 100 cover letters from Study 2, and rated them on the 5 principles. The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001, difference in rs, p = .56,). These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 3, and 4, respectively).
To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters. We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S7, S20, and S29), suggesting that our effects are not explainable by same-model bias.
Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 3, and 4, respectively).
@@
For Study 1, we conducted one-sample t-tests comparing mean responses to the scale midpoint.
For Studies 2–4, as per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness). We used logistic regression to predict whether participants chose to see optional test feedback from condition, controlling for pretest score and baseline characteristics. Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1. When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (benjamini1995controlling).
@@ D Results Study 3
- D1 Randomization, Balance, and Missingness
- D2 Details on professional editors
- D3 Details on participants’ search behavior
- D4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
- D5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
- D6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
- D7 Results were not moderated by individual differences
- D8 Pairwise comparisons
- D9 Forecasting and Willingness to Pay
@@ E Results Study 4
- E2 AI examples improve writing skill
- E3 Seeing AI examples was less effortful
- E4 Seeing an AI example did not discourage motivation for future learning
- E5 The benefits of seeing an AI example were just as large a day later
- E6 Seeing AI examples was equally effective across subgroups
- E7 Pairwise Comparisons
@@ Appendix A Additional methods
We used a randomly selected sample of 100 cover letters from Study 2 to validate the AI ratings. Two trained raters independently rated each of the 5 principles. The inter-rater correlation was r = .74, p < .001. The human raters correlated with the AI-generated ratings satisfactorily (r = .70, p < .001). The correlation indicating AI-human agreement was not smaller than the interrater agreement correlation (difference in rs = .04, p = .56), suggesting that AI ratings achieved similar levels of consistency as human raters. See Figure S1.
@@ Appendix B Results Study 1
We partnered with the Gallup organization to conduct a nationally representative survey of young adults (18-28 years old, N = 2,472). All analyses use survey weights to ensure representativeness of the U.S. population.
B1 Usage
We queried participants about the frequency with which they used AI for different purposes. Table S2 shows the percentage of participants who engaged in these uses at least once in the past month.
| Use Case | Percentage |
| Google replacement | 64.7 |
| Work | 51.5 |
| Writing | 45.9 |
| Image generation | 39.3 |
| Life advice | 32.2 |
| Math | 29.8 |
| Friend/companionship | 22.9 |
| Cheating | 16.4 |
| Romantic advice | 9.9 |
B2 Attitudes
Participants answered four items assessing their beliefs about AI and its effects on human capabilities: “AI makes people less smart”, “AI reduces people’s ability to learn”, “AI makes people more smart” (reverse-coded), and “AI makes people lazier”.
The four items showed good internal consistency (Cronbach’s = 0.84; weighted = 0.80).
All four items showed significantly more negative attitudes than the scale midpoint (all ps < .001). Specifically, 63% agreed that AI makes people less smart, 62% agreed that AI reduces the ability to learn, and 79% agreed that AI makes people lazier. Overall, 73% of respondents scored above the midpoint on a composite variable composed of these 4 items, indicating predominantly negative attitudes. See Table S3
| Item | M | SE | df | Cohen’s | ||
| “AI makes people less smart” | 3.93 | 0.04 | 10.3 | 2636 | 0.30 | |
| “AI reduces ability to learn” | 3.99 | 0.04 | 11.8 | 2636 | 0.34 | |
| “AI makes people more smart” (rev.) | 2.80 | 0.04 | -18.9 | 2636 | -0.55 | |
| “AI makes people lazier” | 4.37 | 0.04 | 22.6 | 2636 | 0.66 | |
| AI Impact Score (composite) | 4.12 | 0.03 | 20.4 | 2636 | 0.57 |
There was no significant relationship between age and the composite measure of AI Impact Score ( = 0.005, p = 0.507).